OpenStudy (anonymous):

If C dollars are into an account paying r percent annual interest, the amount after x years is A=C(1+r)^x. Approximate the amount in the account after x years, when C=$100, R=25%, and X=20 years.

jimthompson5910 (jim_thompson5910):

Plug in the given info to get A=C(1+r)^x A=100(1+0.25)^20 A = ???

OpenStudy (anonymous):

I got to that part but got stuck and unable to get it figured out from there once I got to A=(100(1.25)^20 wasn't able to get my calculator to work right

jimthompson5910 (jim_thompson5910):

what calculator do you have

jimthompson5910 (jim_thompson5910):

you should have 100(1.25)^20 and not (100(1.25)^20

jimthompson5910 (jim_thompson5910):

notice there's one extra ( in the second expression

OpenStudy (anonymous):

I am using a TI-30xA

jimthompson5910 (jim_thompson5910):

ok I've never used that calculator at all I recommend using either google or wolfram alpha type "100(1.25)^20" into either without quotes and tell me what you get

jimthompson5910 (jim_thompson5910):

this is what I mean with google https://www.google.com/search?q=100%281.25%29^20

OpenStudy (anonymous):

8673.61737988 this is the answer I got

jimthompson5910 (jim_thompson5910):

google acts like a calculator for searches like that

jimthompson5910 (jim_thompson5910):

same here, so $8,673.62 is the answer

OpenStudy (anonymous):

k thank you for your help I appreciate it

jimthompson5910 (jim_thompson5910):

you're welcome